OpenStudy (sidsiddhartha):

what is the bilateral laplace transform of the function \[\Huge x(t)=e^t* \frac{ d }{ dt }[e^{-2t}u(-t)]\]

OpenStudy (kainui):

What's a bilateral laplace transform?

OpenStudy (sidsiddhartha):

yeah i am also unsure of it. is it just simple laplace transform?

OpenStudy (kainui):

I've never heard of a "bilateral" one. I was hoping you had a definition or something. We can try googling it.

OpenStudy (anonymous):

It just means the limits are \(-\infty\to \infty\).

OpenStudy (sidsiddhartha):

wiki says http://en.wikipedia.org/wiki/Two-sided_Laplace_transform so just the limits are -inf to inf

OpenStudy (sidsiddhartha):

yeah @wio

OpenStudy (anonymous):

is \(* \equiv \cdot\)?

OpenStudy (sidsiddhartha):

yeah .

OpenStudy (anonymous):

\[ \int\limits_{-\infty}^{\infty} e^{(1-s)t}\frac{d[e^{-2t}u(-t)]}{dt} \]

OpenStudy (anonymous):

Any idea what \(u\) is?

OpenStudy (anonymous):

Is it heaviside?

OpenStudy (sidsiddhartha):

yeah u(t) means unit step function or heaviside function

OpenStudy (anonymous):

It splits into two integrals.

OpenStudy (sidsiddhartha):

like \[\int\limits_{-\infty}^{0^-}+\int\limits_{0^-}^{\infty}\]

OpenStudy (anonymous):

Basically.

OpenStudy (sidsiddhartha):

yeah then first one will be zero is'nt it?

OpenStudy (anonymous):

And fortunately, one of them should go to \(0\), right?

OpenStudy (sidsiddhartha):

yeah ok now its just simple laplace

OpenStudy (anonymous):

This heaviside function is flipped though, so I think the negative side stays.

OpenStudy (sidsiddhartha):

how? can we really write that

OpenStudy (anonymous):

\[ u(t) = \begin{cases} 0&t<0\\ 1&t\geq 0\\ \end{cases} \]Now \[ 1-u(t) =1- \begin{cases} 0&t<0\\ 1&t\geq 0\\ \end{cases} = \begin{cases} 1&t<0\\ 0&t\geq 0\\ \end{cases} = u(-t) \]I suppose with the exception being \(0\)?

OpenStudy (anonymous):

Though in an integral, you can take out any point and it still works...

OpenStudy (sidsiddhartha):

okay understood :) i think now i can do it

OpenStudy (sidsiddhartha):

but one thing when we are writing u(-t)=1-u(t) ,does the region of convergence(ROC) will remain unchanged ?

OpenStudy (anonymous):

By the way, is \[ \frac{d}{dt}u(t)f(t) = u(t)\frac{d}{dt}f(t) \]Since \(u(t)\) is two constants?

OpenStudy (anonymous):

Eh... now that I think about it... it is a bit risky to use that since they aren't equal at \(0\).

OpenStudy (sidsiddhartha):

yeah and for the derivative part i think i will use \[\large \frac{ d }{ dt }x(t)=s.x(s)\]

OpenStudy (anonymous):

Interestingly enough: \[ \lim_{t\to 0^+}u(-t) = 0 = \lim_{t\to 0^+}1-u(t) \]And \[ \lim_{t\to 0^-}u(-t) = 1 =\lim_{t\to 0^-}1-u(t) \]

OpenStudy (anonymous):

The only issue really is that \(u(-0) \neq 1-u(0)\)

OpenStudy (anonymous):

Originally, \(u(0)\) was arbitrary or undefined anyway.

OpenStudy (sidsiddhartha):

yeah i think u(0) will result a impulse

OpenStudy (anonymous):

It's almost sort of a meaningless value.

OpenStudy (anonymous):

\[\begin{split} \int\limits_{-\infty}^{\infty} e^{(1-s)t}\frac{d[e^{-2t}u(-t)]}{dt} &=\int\limits_{-\infty}^{0} e^{(1-s)t}\frac{d[e^{-2t}(1)]}{dt}+\int\limits_{0}^{\infty} e^{(1-s)t}\frac{d[e^{-2t}(0)]}{dt}\\ &=\int\limits_{-\infty}^{0} e^{(1-s)t}\frac{d[e^{-2t}]}{dt}\\ &=\int\limits_{-\infty}^{0} e^{(1-s)t}(-2e^{-2t})dt\\ \end{split} \]\[ \]

OpenStudy (anonymous):

\[ =2\int\limits_0^{\infty}e^{(1-s-2)t}dt \]

OpenStudy (anonymous):

I think you can do the rest... unless I made a mistake?

OpenStudy (sidsiddhartha):

yeah thank u ^_^ i will do it